OpenStudy (anonymous):

What are the transformations of g(x) = loga(3-x) + 1 and t(x) = 2loga 3x

OpenStudy (jdoe0001):

\(\large { g(x) = log_a(3-x) + 1\\ \quad \\\implies \begin{array}{cccllll} g(x) = log_a(&-1x+3)& + 1\\ &\uparrow &\uparrow \\ &\textit{horizontal shift}&\textit{vertical shift}\\ &\frac{3}{-1}&+1 \end{array} \\ \quad \\ \quad \\ \begin{array}{llll} t(x) = &2log_a &3x\\ &\uparrow &\uparrow \\ \end{array}\\\qquad \textit{vertical stretch} }\)

OpenStudy (anonymous):

so for t(x) = 2loga 3x. Its a vertical stretch of 3 or two? a little confused

OpenStudy (jdoe0001):

http://fooplot.com/#W3sidHlwZSI6MCwiZXEiOiJsb2coeCkiLCJjb2xvciI6IiMwMDAwMDAifSx7InR5cGUiOjAsImVxIjoiMmxvZyh4KSIsImNvbG9yIjoiI0U2MTAxMCJ9LHsidHlwZSI6MCwiZXEiOiIybG9nKDN4KSIsImNvbG9yIjoiI0U4MTJDRiJ9LHsidHlwZSI6MTAwMH1d <--- they both create a stretch on the graph of the function this is logarithmic function.. so it goes the the right side so the factors will expand the value for "y" thus it'll mean a horizontal "expansion"